# Learning-Augmented Priority Queues

### v1

@@ 6 Experiments
In this section, we empirically evaluate the performance of our learning-augmented priority queue (LAPQ) by comparing it with Binary and Fibonacci heaps. We use two standard benchmarks for this evaluation: sorting and Dijkstra’s algorithm. We also compare our results with those of Bai and Coester [2023] for sorting. For Dijkstra’s algorithm, we assess performance on both real city maps and synthetic random graphs. In all the experiments, each data point represents the average result from 30 independent runs.
@@ 6.1 Sorting
Figures 2 and 3 show the obtained results respectively in the class and the decay setting for $n\in\{10^{3},10^{4},10^{4}\}$. In the class setting with offline predictions, the LAPQ slightly outperforms the Double-Hoover and Displacement sort algorithms of Bai and Coester [2023], which were shown to outperform classical sorting algorithms. In the decay setting, the LAPQ matches the performance of the Displacement sort, but is slightly outperformed by the Double-Hoover sort. With online predictions, although the problem is harder, LAPQ’s performance remains comparable to the previous algorithms. In both settings, the LAPQ with offline predictions, online predictions, and dirty comparisons all yield better performance than binary or Fibonacci heaps, even with predictions that are not highly accurate.
@@ 6.2 Dijkstra’s algorithm
For this, we use Poisson Voronoi Tessellations (PVT), which are a commonly used random graph model for street systems Gloaguen et al. [2006]; Gloaguen and Cali [2018]; Benomar et al. [2022a, b].
@@
We present the results in the class and the decay settings respectively in Figures 8 and 7. Similar to previous experiments with Dijkstra on city maps, these figures illustrate how the number of comparisons decreases when the LAPQ is augmented with node rank predictions or with the corresponding dirty comparator. We compare them with the number of comparisons induced by using a binary or Fibonacci heap, as well as with the number of comparisons of the LAPQ augmented with key rank predictions.

### v2

@@ 6 Experiments
In this section, we empirically evaluate the performance of our learning-augmented priority queue (LAPQ) by comparing it with Binary and Fibonacci heaps. We use two standard benchmarks for this evaluation: sorting and Dijkstra’s algorithm. We also compare our results with those of Bai and Coester [2023] for sorting. For Dijkstra’s algorithm, we assess performance on both real city maps and synthetic random graphs. In all the experiments, each data point represents the average result from 30 independent runs. The code used for conducting the experiments is available at github.com/Ziyad-Benomar/Learning-augmented-priority-queues.
@@ 6.1 Sorting
Figures 2 and 3 show the obtained results respectively in the class and the decay setting for $n\in\{10^{3},10^{4},10^{5}\}$. In the class setting with offline predictions, the LAPQ slightly outperforms the Double-Hoover and Displacement sort algorithms of Bai and Coester [2023], which were shown to outperform classical sorting algorithms. In the decay setting, the LAPQ matches the performance of the Displacement sort, but is slightly outperformed by the Double-Hoover sort. With online predictions, although the problem is harder, LAPQ’s performance remains comparable to the previous algorithms. In both settings, the LAPQ with offline predictions, online predictions, and dirty comparisons all yield better performance than binary or Fibonacci heaps, even with predictions that are not highly accurate.
@@ 6.2 Dijkstra’s algorithm
For this, we use Poisson Voronoi Tessellations (PVT), which are a commonly used random graph model for street systems Gloaguen et al. [2006]; Gloaguen and Cali [2018]; Benomar et al. [2022, ].
@@
We present the results in the class and the decay settings respectively in Figures 7 and 8. Similar to previous experiments with Dijkstra on city maps, these figures illustrate how the number of comparisons decreases when the LAPQ is augmented with node rank predictions or with the corresponding dirty comparator. We compare them with the number of comparisons induced by using a binary or Fibonacci heap, as well as with the number of comparisons of the LAPQ augmented with key rank predictions.